50. In Parts (a)-(b), you are given a pair of ordered bases B and B for R2. Find the change of coordinate matrix that changePlease finish all the problems. I will really appreciate it.
Answer #1

(5 @ 8= {(1,3), (2, 517 and ß = {(1,0), 10,157 T[(1,0)= a (1,3) + B (2,5) T[(1,01] = 12,34 ) + (28, 5B) (10)] = (x+2/3 , 34+

8 B = {1,01, (0,177 and ß = {(1,3), 12,51 4. T[ 11,37] = (do T(1,3)] = (1,3)] = 2(1,0)7 810,1) (0,0) +10.ß) T(1,3)] = (248) 2

: 20-33=7 - T(Cd) =-3 (1.1) +4(10) Ted) = (-3,-3) +14,0) Tead) = (1-3) B = 4 (1,2), (1, -3) Q+Bz) - 0 2x & (1) - Can ② 24+2